HYPERFINE SPLITTING AND ROTATIONAL ANALYSIS OF THE DIATOMIC MOLECULE ZINC MONO-SULFIDE, ZnS.^{a}

\underline{DANIEL J. FROHMAN}, G. S. GRUBBS II, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Hall-Atwater Laboratories, 52 Lawn Avenue, Middletown, CT 06459-0180.

67Zn hyperfine structure has been observed in the diatomic molecule ZnS in the microwave (6-26 GHz) region. The molecule was synthesized by the use of a newly constructed laser ablation source based on the design of Walker and Gerry.^{b} Previous rotational studies of this molecule have been performed in the millimeter-wave region (370-471 GHz range).^{c} Rotational analyses, including the nuclear electric quadrupole coupling constant, will be discussed and compared with the literature.

aSupport from CHE-1011214

bK. A. Walker and M. C. L. Gerry, J. Mol. Spectrosc., \underline{182} (1997), 178

cL. N. Zack and L. M. Ziurys, J. Mol. Spectrosc., \underline{257} (2009), 213-216